OpenStudy (sarahezzmck):

Can someone help me with Solving Exponential Equations with unequal bases?

OpenStudy (anonymous):

Sure

OpenStudy (sarahezzmck):

ok i have to do my equation on inetrest rates, and the payment of a car. The car payment is 5,990 and the interest rate is 4.84 for 3 months. Im not sure how I can set this up or how to solve it :/

OpenStudy (anonymous):

What are you trying to find? I.e., what is the problem asking for? You stated that you have to do a problem on interest rates and payment, but you already told me the interest rate (4.94%) and payment (5,990).

OpenStudy (sarahezzmck):

well the question says Using the function A(t)=P(1+ r/n)^nt , create the function that represents your new car loan. The principle will be the price of the vehicle you selected, not how much you are putting down. Im not sure how to create the function

OpenStudy (anonymous):

what is the question

OpenStudy (sarahezzmck):

thats the question the one that i just posted. thats what the question says, I copied and pasted it word for word. My price for the vehicle is 5,990 and the interest rate that i looked up is 4.84% @Rolltideroll

OpenStudy (anonymous):

Alright, so that is the compound interest formula. P is the principle, r is the interest rate, t is the time that the money is in the account, and n is the number of times the money is compounded per year.

OpenStudy (anonymous):

What is the price of the vehicle - the principle? is it 5900? or is that "the payment" as you typed?

OpenStudy (anonymous):

I=(Pxrxi)/12 = (5990x0.0484x3)/12 = the answer

OpenStudy (sarahezzmck):

its the principle

OpenStudy (sarahezzmck):

lol, thanks Moe.

OpenStudy (anonymous):

ur wlcm

OpenStudy (sarahezzmck):

i still need help tho guys, that wasnt the end of it. If thats alright :/

OpenStudy (anonymous):

continue plz

OpenStudy (sarahezzmck):

It says "Use your function to determine how much you will owe the bank in five years. This is assuming you are not paying down your loan, so do not get scared! Show all of your work." how do I figure out how much I will owe the bamk in 5 years? Do I multiply the answer?

OpenStudy (anonymous):

i dont know i forgot how to do this

OpenStudy (sarahezzmck):

lol is ok Alex thanks for trying though :)

OpenStudy (anonymous):

no problem what type of math is this

OpenStudy (sarahezzmck):

Algebra II

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Do you know how many times it's compounded per year

OpenStudy (sarahezzmck):

um no, wouldnt it be 12 or 4.84? Or the answer to the last one which was 7.2

OpenStudy (anonymous):

Using the formula you gave I get: A(t) = 5990 [ 1 + (0.0494 / n)]^[n(5)] where n is how many times it's compounded annually

OpenStudy (sarahezzmck):

So how do you get n?

OpenStudy (anonymous):

It should be given for you in the problem. if it's not, then I assume n would be 1

OpenStudy (sarahezzmck):

oh ok, bc it's not given at all.

OpenStudy (anonymous):

Some problems would say "... blah blah blah.. compounded quarterly..." which would make n = 4.... but if your problem says nothing, I think it is 1

OpenStudy (sarahezzmck):

ok, yes it doens't say anything about the annual compound.

OpenStudy (anonymous):

is it a simple or a compound interest?

OpenStudy (sarahezzmck):

um, i think its a compound interest since it has to do with interest rates and such.. :/

OpenStudy (sarahezzmck):

i can post the whole page if you guys would like

OpenStudy (anonymous):

"well the question says Using the function A(t)=P(1+ r/n)^nt , create the function that represents your new car loan. T" = compound interest

OpenStudy (sarahezzmck):

oh ok, so yeah compound interest sorry!!!

OpenStudy (anonymous):

:) Well I think we're done here. All finished right?

OpenStudy (sarahezzmck):

lol noo, its a really long page :/

OpenStudy (anonymous):

lol

OpenStudy (sarahezzmck):

im sorry Im a pain! I just really dont understand math

OpenStudy (anonymous):

can u attach the file?

OpenStudy (anonymous):

Nobody does, it requires time and effort. Some people are great and pick everything up very fast, others require more effort. I'm pretty average but I made it to Calculus 3 and other advanced topics. You can do it!

OpenStudy (sarahezzmck):

yes

OpenStudy (anonymous):

(If I can do it, anyone can)

OpenStudy (sarahezzmck):

lol thanks :)

OpenStudy (anonymous):

oh cool well i have to step out but message me if nobody helped you and when I get back i can help with your assignment

OpenStudy (anonymous):

30 min or so

OpenStudy (sarahezzmck):

ok thanks so much btw!

OpenStudy (sarahezzmck):

@ranga are you here to help me too? I hear ur like a genius!!!

OpenStudy (anonymous):

hold on let me see if one of my other friends can help you with this he is very good at math

OpenStudy (sarahezzmck):

omg thnks Alex!!!!!!!!!!

OpenStudy (anonymous):

@amoodarya can you help Sahar with her question

OpenStudy (anonymous):

no prob

OpenStudy (sarahezzmck):

:D

OpenStudy (anonymous):

@amoodarya

OpenStudy (ranga):

Car loans tend be compounded monthly. So I would choose n = 12. P = $5,990 r = 4.84% = 0.0484 n = 12 (how many times the interest is compounded per year) A(t) = P * (1 + r/n)^(nt) = 5900 * (1 + 0.0484 / 12)^(12*t) A(t) = 5900((1.0040333)^(12t) is the function.

OpenStudy (anonymous):

bump your question

OpenStudy (sarahezzmck):

OMG thank you @ranga and i cant Alex

OpenStudy (anonymous):

oh ok

OpenStudy (ranga):

The 4.84% interest rate that you looked up, is that 4.84% APR? (APR-Annual Percentage Interest)

OpenStudy (sarahezzmck):

um, im not sure, i went on google and i just looked up, "Current Auto Loan Interest rates" and it brought me to this website: http://www.bankrate.com/finance/auto/current-interest-rates.aspx and I chose the January 22nd, 2014

OpenStudy (ranga):

They don't explicitly say so we will have to assume it is annual interest rate in which case the A9t) I gave above is correct.

OpenStudy (anonymous):

hey become a fan of @amoodarya

OpenStudy (sarahezzmck):

ok :) and then my problem says, use the function to determine how much you will owe in 5 years, This is assuming you are not paying your down loan. How do I figure that out (i did alex)

OpenStudy (anonymous):

ok send him the question and let him look at it

OpenStudy (sarahezzmck):

im getting help now, i'll send another question to him if I have one Alex

OpenStudy (ranga):

So for the first part the function is: A(t) = 5900(1.0040333)^(12t) For the part that asks how much you will owe the bank in 5 years if no monthly payments are made, just plug in t = 5 in the formula and calculate A. A(t) = 5900 * (1.0040333)^(12t) Put t = 5 A = 5900 * (1.0040333)^(12 * 5) A = 5900 * (1.0040333)^(60) A = $7511.71 In 5 years, if no monthly payments is made, you will owe $7511.71 to the bank on a $5,900 loan.

OpenStudy (sarahezzmck):

Thank you so much!! Can u help me with more please???

OpenStudy (ranga):

You are welcome. I can try one more. But post it as a new question because this is getting too long.

OpenStudy (sarahezzmck):

okeydoke :)